40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9228     Entry time: Wed Oct 9 22:58:34 2013
Author: Manasa 
Type: Update 
Category: Green Locking 
Subject: ALS stabilization 

After Jenne and Masayuki told that they were not able to stabilize the ALS for either arms yesterday, I looked into things with the ALS servo.

I had trouble initially trying to even stabilize the loop for a few minutes. So I measured the OLTF of the phase tracker loop and the ALS X arm servo. I changed phase tracker gain to 125 and that rendered UGF of 2KHz and phase margin of 45 degrees for the phase tracker loop.

The ALS servo gain was set such that UGF was 125Hz and phase margin 38 degrees (attached is the transfer function measurement for the servo).

I could stabilize the arm to ~500 Hz/rtHz (rms), which is twice that of what we had while we did the (PRMI+1arm ALS).

But ALS was still not stable long enough with the higher rms to even allow a cavity scan to find IR resonance. I suspect the problem to now lie with the PDH loop. We should be looking to stabilize the PDH for green if we need a stable ALS.

Attachment 1: ALS_XARM_OLTF.pdf  11 kB  | Hide | Hide all
ALS_XARM_OLTF.pdf ALS_XARM_OLTF.pdf ALS_XARM_OLTF.pdf
ELOG V3.1.3-